OpenStudy (lexy148):

how do you find the degree of a right triangle with cosine of 39 and 44

OpenStudy (amistre64):

needs a bit of clarification

OpenStudy (anonymous):

i'll say

OpenStudy (amistre64):

what do 39 and 44 tell us?

OpenStudy (lexy148):



OpenStudy (amistre64):

well the marked angle is primed for sin, not cos .... other than that

OpenStudy (amistre64):

for practice, can you tell me the definitions of sin, cos in terms of their side ratios?

OpenStudy (lexy148):

sin=opp/hyp cos=adj/hyp and tan=opp/adj

OpenStudy (amistre64):

so in this case, we have an angle, and the sides over from it, and hypot

OpenStudy (amistre64):

which trig function would we use

OpenStudy (amistre64):